### v2

@@ B.4 Refinement Feature Injection for Enhanced Consistency
B.5 Implementation details
Anchor Videos: Similar to ConsiStory, we use two anchor videos that share all features between themselves. Further implementation details are provided in Appendix B.15
B.6 Consistent Video Generation - Comparisons to Baselines
@@
Figure 9: Qualitative Comparisons. (click-to-view-online) The first frame of each video shot is displayed along with a spatiotemporal y–t slice to visualize motion. Ours (top row) shows improved character consistency across shots while maintaining natural motion. VideoCrafter2 (row 2) is the vanilla model, showing diverse motion but inconsistent characters. Tokenflow-Encoder (row 3) preserves original motion but struggles with character consistency and introduces coloring artifacts. ConsiS Im2Vid (bottom row) fails to maintain consistency and exhibits limited motion adherence to prompts. See more examples in Fig. 14.
@@
B.7 Evaluation baselines
We compare our method to several baselines: (1) VideoCrafter2: A baseline “vanilla” text-to-video model (Chen et al., 2024), without adaptations. VideoCrafter2 is a public SoTA video model (Huang et al., 2024). (2) Tokenflow-Encoder: A combination of TokenFlow Geyer et al. (2023) with IP-Adapter, a Personalization-Based Encoder (Ye et al., 2023). We personalize TokenFlow by conditioning the IP-Adapter on the first frame of one video generated by the vanilla model. For IP-Adapter we use a high-scale hyper-parameter to push the model toward stronger consistency. (3) ConsiS Im2Vid: A combination of SoTA image-consistency approach (Tewel et al., 2024), with a subsequent Image-to-Video variant of VideoCrafter (Chen et al., 2024). First, we generate a set of consistent reference images. Then, we use them as inputs to an Image-to-Video model. We chose VideoCrafter, as it is a public image-to-video model that has an overall quality equivalent to that of the text-to-video VideoCrafter2 model according to the VBench benchmark Huang et al. (2024). (4) VSTAR: A method for generating a long video with dynamic evolution (Li et al., 2024b). We directly provide the multiple prompts and sample 16 frames per prompt, then splitting the result into individual shots. (5) Turbo-V2: A recent state-of-the-art text-to-video model Li et al. (2024a) that we use to demonstrate our method’s adaptability to other architectures.
B.8 Qualitative Results
@@
Figure 10: Ablation Study on ConsiStory Components for Video Generation. (click-to-view-online) “Ours” (top row) demonstrates improved motion richness and identity preservation. VideoCrafter2 (second row) shows diverse motion but inconsistent characters. “ConsiS” (third row), a naive ConsiStory implementation, shows impaired identity and motion artifacts. “ConsiS +Uncond” (fourth row) adds feature injection to unconditional denoising, resolving motion artifacts but reducing motion magnitude and compromising identity. “Q ConsiS” (fifth row) couples each frame with a single frame in an anchor video, allowing some natural motion, although partially synchronized, with improved identity. Our method achieves the best balance of motion quality and identity.
@@ B.10 Additional Results
Figure 13: T2V-Turbo-V2: (click-to-view-online) Video Storyboarding can be applied to T2V-Turbo-V2 Li et al. (2024a), a recent state-of-the-art video model, that exhibits significantly better motion.
@@
Figure 14: Additional Qualitative Comparisons , including VSTAR: (click-to-view-online) Our method generates consistent subjects while preserving diverse and natural motions across scenarios.
@@
B.11 Framewise Subject-Driven Self-Attention - Implementation Details
@@
B.12 Flow-based Q components injection - Formal Definition
@@
B.13 Benchmark Dataset Construction:
@@ B.15 Implementation Details
B.16 User Study Protocol

### v3

@@ B.4 Refinement Feature Injection for Enhanced Consistency
B.5 Consistent Video Generation - Comparisons to Baselines
@@
Figure 11: Qualitative Comparisons. The first frame of each video shot is displayed along with a spatiotemporal y–t slice to visualize motion. Ours (top row) shows improved character consistency across shots while maintaining natural motion. VideoCrafter2 (row 2) is the vanilla model, showing diverse motion but inconsistent characters. Tokenflow-Encoder (row 3) preserves original motion but struggles with character consistency and introduces coloring artifacts. ConsiS Im2Vid (bottom row) fails to maintain consistency and exhibits limited motion adherence to prompts. See more examples in Fig. 16.
@@
B.6 Evaluation baselines
We compare our method to several baselines: (1) VideoCrafter2: A baseline “vanilla” text-to-video model, without adaptations. VideoCrafter2 is a public SoTA video model (Huang et al., 2024). (2) Tokenflow-Encoder: A combination of TokenFlow Geyer et al. (2023) with IP-Adapter, a Personalization-Based Encoder (Ye et al., 2023). We personalize TokenFlow by conditioning the IP-Adapter on the first frame of one video generated by the vanilla model. For IP-Adapter we use a high-scale hyper-parameter to push the model toward stronger consistency. (3) ConsiS Im2Vid: A combination of SoTA image-consistency approach (Tewel et al., 2024), with a subsequent Image-to-Video variant of VideoCrafter. First, we generate a set of consistent reference images. Then, we use them as inputs to an Image-to-Video model. We chose VideoCrafter, as it is a public image-to-video model that has an overall quality equivalent to that of the text-to-video VideoCrafter2 model according to the VBench benchmark Huang et al. (2024). (4) VSTAR: A method for generating a long video with dynamic evolution (Li et al., 2024b). We directly provide the multiple prompts and sample 16 frames per prompt, then splitting the result into individual shots. (5) Turbo-V2: A recent state-of-the-art text-to-video model Li et al. (2024a) that we use to demonstrate our method’s adaptability to other architectures.
B.7 Qualitative Results
@@
Figure 12: Ablation Study on ConsiStory Components for Video Generation. “Ours” (top row) demonstrates improved motion richness and identity preservation. VideoCrafter2 (second row) shows diverse motion but inconsistent characters. “ConsiS” (third row), a naive ConsiStory implementation, shows impaired identity and motion artifacts. “ConsiS +Uncond” (fourth row) adds feature injection to unconditional denoising, resolving motion artifacts but reducing motion magnitude and compromising identity. “Q ConsiS” (fifth row) couples each frame with a single frame in an anchor video, allowing some natural motion, although partially synchronized, with improved identity. Our method achieves the best balance of motion quality and identity.
@@ B.9 Additional Results
Figure 15: T2V-Turbo-V2: Video Storyboarding can be applied to T2V-Turbo-V2 Li et al. (2024a), a recent state-of-the-art video model, that exhibits significantly better motion.
@@
Figure 16: Additional Qualitative Comparisons , including VSTAR: Our method generates consistent subjects while preserving diverse and natural motions across scenarios.
@@
B.10 Framewise Subject-Driven Self-Attention - Implementation Details
@@
B.11 Flow-based Q components injection - Formal Definition
@@
B.12 Benchmark Dataset Construction:
@@ B.14 Implementation Details
B.15 User Study Protocol